OpenStudy (anonymous):

Solve: 2(x - 3) + 2(x - 7) = 40

OpenStudy (anonymous):

so distribute 2 iinto the first equation

OpenStudy (anonymous):

you get 2x-6

OpenStudy (anonymous):

2x-6

OpenStudy (anonymous):

do the same thing for the second one

OpenStudy (anonymous):

2x-14

OpenStudy (anonymous):

yes so you get 2x-6+2x-14=40

OpenStudy (anonymous):

then what

OpenStudy (anonymous):

now add the x together and the numbers together

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

4x-20+40?

OpenStudy (anonymous):

4x-20=40

OpenStudy (anonymous):

then move the 20 to the other side

OpenStudy (anonymous):

4x=40-20

OpenStudy (anonymous):

it becomes 40+20 not - becuase u nmove it to the other side

OpenStudy (anonymous):

so 4x= 40+20

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so 4x=60

OpenStudy (anonymous):

so x =?

OpenStudy (anonymous):

15

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

could i get medal pls

OpenStudy (anonymous):

i gave you one

OpenStudy (anonymous):

thanks